OpenStudy (nuccioreggie):

At Priscilla's Purse Warehouse, it takes fraction 2 over 3 of a day to complete fraction 1 over 9 of an order of purses. At this rate, how long will it take to complete the entire order of purses? Mixed number 3 and 2 over 3 days 5 days 6 days Mixed number 13 and 1 over 2 days

OpenStudy (anonymous):

multiply 2/3 by the 9 days it will take

OpenStudy (nuccioreggie):

@jakashaka123

OpenStudy (anonymous):

yeah ?

OpenStudy (anonymous):

\(\frac{2}{3}\) days to complete \(\frac{1}{9}\) of the job?

OpenStudy (nuccioreggie):

Ok

OpenStudy (anonymous):

\[\frac{2}{3}\times 9\]

OpenStudy (anonymous):

6days...

OpenStudy (anonymous):

is six a choice?

OpenStudy (nuccioreggie):

Yea i think its c

OpenStudy (anonymous):

yes 6 is the right answer lol :) I think it is c in fact